in another population, you genotype 1000 turtles, and observe that 700 are TT100 are tt and 200 are Tt

What are the observed values of p and q in this population?
Answer #1

Allele frequency = (2*number of homozygous individuals + number of hetrozygous individuals)/2*total number of individuals.

We multiply total number of individuals by 2 as their are two alleles for a gene in a diploid organism, multiply number of homozygous individual by two as there are two copies of same allele, and simply add hetrozygous as there is only one copy of allele that's allele frequency needed to calculate.

p represents the allele frequency of dominant allele

p =( 2*700+200)/2000

1600/2000 = 0.8

q represents the allele frequency of recessive allele

q = 1-0.8 = 0.2 ; using equation p+q = 1

(= 2*100+200/2000

= 400/2000

= 0.2)